OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 4 over 5(4/5), what is the equation of the line?

OpenStudy (anonymous):

@gfeder

OpenStudy (anonymous):

y=x+1

OpenStudy (anonymous):

a.4y − 5x = 5 b.y − 4x = 5 c.5y + 4x = 5 d.5y − 4x = 5 these are the choices

OpenStudy (mrnood):

write the equation as y=mx +b where m is the slope (4/5) and b is the y value where the line crosses y axis. You then need to compare that equation with the 4 options (you may need to re-arrange them to look the same (y=mx+b)